### v1

# Enhancing Vision-Language Models with Scene Graphs for Traffic Accident Understanding

Abstract
Recognizing a traffic accident is an essential part of any autonomous driving or road monitoring system. An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from reoccurring. The task of being able to classify a traffic scene as a specific type of accident is the focus of this work. We approach the problem by likening a traffic scene to a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges. This representation of an accident can be referred to as a scene graph, and is used as input for an accident classifier. Better results can be obtained with a classifier that fuses the scene graph input with representations from vision and language. This work introduces a multi-stage, multimodal pipeline to pre-process videos of traffic accidents, encode them as scene graphs, and align this representation with vision and language modalities for accident classification. When trained on 4 classes, our method achieves a balanced accuracy score of 57.77% on an (unbalanced) subset of the popular Detection of Traffic Anomaly (DoTA) benchmark, representing an increase of close to 5 percentage points from the case where scene graph information is not taken into account.
Index Terms:
Neuro-symbolism, Vision-Language Models, Foundation Models, Autonomous Driving, Perception
I Introduction
The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicle (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022). A major aspect of this task is to efficiently and accurately recognize different types of traffic accidents, with the ultimate goal of preventing them. In this work, we approach this challenge by modeling traffic scenes using scene graphs. To improve classification performance, the graph representation is further aligned with representations from the vision and language modalities within contrastively-trained foundation models. Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline to pre-process videos, encode them as scene graphs, and align these representations with vision and language modalities to classify traffic accidents.
In this paper, we propose Scene-Traffic-Graph Inference(STGi) as a unified system for accident classification. First, we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al. (2021). Then, we use a scene graph encoder (SGE) from the same authors Malawade et al. (2022) to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al. (2021) and X-CLIP Ni et al. (2022), respectively. X-CLIP directly expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames. In essence, scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and are all used together to classify traffic accident scenes.
A summary of contributions is as follows:
  • We introduce Scene-Traffic-Graph Inference (STGi), a novel method for traffic accident classification which leverages scene graphs to capture the essential features of a traffic accident.
  • We demonstrate that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
  • Experiments in this work demonstrate that aligning the scene graph modality with vision and language give similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
Figure 1: An overview of the STGi Architecture. Beginning with video and text inputs, video frames are sampled and used to generate scene graphs. Then, alignment training is performed on the three encoders before a prediction head is used to classify the accident.
II Related Work
II-A Scene Graphs for Representing Traffic Scenes
Several works have taken the approach of modeling traffic scenarios using scene graphs (SGs). Yu et al. (2020) propose one such representation that forms the basis of the SGs produced in our work. They define a graph structure with nodes representing entities such as vehicles and pedestrians while edges represent relations between the objects and are labeled by distance category (e.g., visible, very near) and orientation (e.g., in front of, beside). Zipfl and Zöllner (2021) also model these objects in a traffic scene, further accounting for relative speeds of objects in the graph edges. Guo et al. (2023) extracts similar semantic information from a traffic scene, but the focus is on generating a knowledge graph, accounting for added traffic information present in a scene such as road signs and lanes. Qasemi et al. (2023) also use a knowledge graph to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents. Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a knowledge graph. The use of a knowledge graph in this case is helpful in contextualizing the traffic scene in a broader traffic monitoring ontology and for answering general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene. Instead, our method is better suited for the task of accident classification by leveraging scene graphs to capture relationships among objects that may be involved in a traffic accident in a given video example. Combining the use of scene graphs with knowledge graphs to model traffic settings, Dimasi (2023) makes use of a knowledge graph to boost the performance of the RelTR (Relation Transformer) model Cong et al. (2022), an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
II-B Leveraging Scene Graphs with Neural Networks
There are many possible designs and applications for leveraging the versatile structure of scene graphs to enhance the performance of neural networks Li et al. (2019); Khademi and Schulte (2020); Rana et al. (2023). More generally, Zhang et al. (2019) describe the ability of scene graphs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features. Although their study is not specific to our traffic domain, they claim that scene graphs derived from images can capture the essential visual features and may outperform images for VQA tasks. Many works have focused on automatically generating scene graphs that can easily be encoded into neural networks for various downstream tasks. The rs2v tool Malawade et al. (2021) generates scene graphs for individual frames of traffic videos and is applied to our work. Alternatively, Nag et al. (2023) design a tool specifically for modeling videos with scene graphs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers. However, they train on the Action Genome Ji et al. (2019) dataset and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
II-C Multimodal Alignment with Contrastive Representations
Radford et al. (2021) proposed a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model. From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as aligning audio Wu et al. (2022) and haptic data Tatiya et al. (2024). Similar to our implementation, Koch et al. (2023) use CLIP’s text encoder to contrastively train 3D scene graphs. The work by Huang et al. (2023) aims to enhance CLIP’s performance by injecting scene graph information into the embedding generated by CLIP’s text encoder before contrastive learning is performed between the image and text embedding spaces. To the best of our knowledge, however, our work is the first to treat scene graphs as an additional modality that is grounded to a text and video encoder. There are also many possible strategies for fusing the outputs of a multimodal model before the downstream task at hand, such as early and late stage fusion involving concatenation Sleeman et al. (2022); Pawłowski et al. (2023), merging Sleeman et al. (2022), or sampling from a shared embedding space Gao et al. (2020). We experiment with several of these late fusion techniques on the three modality embeddings after training for alignment.
III Methodology
To approach the problem of traffic accident classification, a multi-stage pipeline was developed. The pipeline, which is represented in Figure 1, can be summarized in four stages and is described in the following sections:
  1. 1.
    Data pre-processing: Sample video frames, generate captions, and use a scene graph generator to produce a set of scene graphs for each traffic accident example.
  2. 2.
    Scene graph encoder pre-training: Pre-train the scene graph encoder on the classification task.
  3. 3.
    Multimodal alignment: Align the scene graph encoder with frozen video and text encoders.
  4. 4.
    Fine tuning for downstream task: Train a classification head on top of the aligned multimodal model.
III-A Data Pre-processing
In the first stage, for extensibility and ease, we leverage and tune existing tools for generating traffic scene graphs that will later be fed into our modeling approach. Various scene graph generators (SGGs) are available for this, although not many fit our specific requirements for a generator that is capable of identifying and encoding features unique to traffic accidents, such as different types of vehicle collisions. Some do work for our needs — in particular, the roadscene2vec (rs2v) tool created by Malawade et al. (2021) which we leverage for generating scene graphs from traffic video frames. To use rs2v, we first sample a fixed amount of frames from each video, and text captions are generated manually for each class. Then, feeding in a series of frames from a traffic scene, the rs2v generator uses an object detector and keeps only the relevant entities relating to traffic (such as road, cars, pedestrians, etc.), filtering out other detected objects. Next, after generating a “bird’s eye view” (BEV) projection of the image and approximating the relative location of each object in this projection, edges are connected between nearby entities in the scene, and vehicles are mapped to lanes. This forms the scene graph for a specific frame from a video sample. The SGG is an important component of this pipeline as it defines the elements in the scene graph modality.
An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used to create edges of varying attributes, such as very near or visible, relating a pair of objects). The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene Malawade et al. (2022). In our case, we iteratively sample and adjust the BEV parameters and proximity thresholds based on the output quality of the SGs produced for various scenes in DoTA. We do this to select one configuration to generate the SGs for our task, although it remains an inherent challenge to generalize the parameter settings for the entire dataset. Some examples of the generated SGs can be seen in Appendix A.
For text, we manually compose captions describing each of the four accident classes (see Appendix B), pairing each caption with videos from its respective class to form the training examples. We also experiment with two sets of captions, however, we focus the scope of this work primarily on the scene graph generation process from video frames rather than caption generation. It should be noted that although these captions are used during training for aligning the SGE, they are not used during inference on the final model because they are not from our dataset and provide a one-to-one mapping directly to the accident classes. Through fine-tuning the classification head using these captions, the model can achieve 100% accuracy this way. Instead, we fine-tune and test the classification head using the same caption as noted in Appendix B for all examples.
III-B Scene Graph Encoder
In stage two, the SGE encodes scene graphs to fixed-length embeddings. To do this, a multirelational graph convolutional network (MRGCN) as employed by Malawade et al. (2022) is used, which includes an attention mechanism along with LSTMs to model the spatial and temporal relations of the scene graphs generated for a given video. The SGE may be pre-trained for the classification task before aligning the scene graphs with the language and vision modalities.
III-C Multimodal Alignment
For stage three, we use the CLIP model Radford et al. (2021), but replace the CLIP image encoder with X-CLIP Ni et al. (2022), a pre-trained video encoder, to accept videos rather than images. We leave CLIP’s text encoder in place. We then freeze the weights from the video and text encoders and align them to the SGE encoder, which takes the new “scene graph modality” as input.
III-D Fine Tuning for Downstream Task
The final stage involves training a classification head that accepts embeddings from the three modalities and outputs an accident class for a traffic scene. This requires the selection of a method to fuse the signal from the three modalities followed by training a small network. For this, we experiment with different late fusion techniques such as taking a weighted linear combination of the modality outputs Kaliciak et al. (2014) and training various MLP classifiers Kiela et al. (2018) with and without activations on top of the concatenated embeddings. Based on these experiments, we choose the approach of concatenating the vectors from the three modalities as described by Sleeman et al. (2022) and, following Wu et al. (2022), train a 2-layer MLP with ReLU activations as our classification head.
IV Experimental Design
IV-A Dataset
Given the nature of our multimodal approach, we require data in the form of videos, text, and scene graphs. Regarding the videos, we focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al. (2022) for this task. DoTA is a dataset consisting of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place. For our use case of this dataset, question one is irrelevant since we only sample frames from the anomalous portion of the video based on the annotations of the DoTA authors. Additionally, question two is less essential for us since we use a SGG, which is meant to extract the necessary objects from video frames and model their relevant spatial relationships for our task. We therefore prioritize providing an answer for the third question, that is, classifying different traffic anomalies. However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to only a subset of four of the available classes.
In addition to the provided video classes, Yao et al. (2022) also divides the videos into two types, ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego. Every video class contains videos of both types. By default, rs2v is designed to only model relationships between the ego vehicle and other objects, not among other objects themselves, so we further restrict our task to the set of ego videos available in DoTA. In total, this leaves us with 4 classes and 2,163 videos. The classes are labeled as follows, describing the type movement of the vehicle as the accident occurs: moving_ahead_or_waiting, oncoming, turning, and lateral.
IV-B Baselines and Metrics
In order to evaluate the usefulness of representing traffic scenes graphically to classify them, we first verify whether a SGE alone is superior to randomly guessing an accident class. We later evaluate the effectiveness of using SGs in a multimodal classifier by comparing the difference in the performance between using just text and video embeddings versus adding in the signal from the SG modality.
For this and other results, the primary metrics we focus on are accuracy and balanced accuracy. Accuracy is simply defined as the proportion of the correctly classified predictions out of the total number of predictions and remains an interpretable metric that is relatively simple to compare across different benchmarks. There is, however, significant imbalance in the dataset, with the turning class appearing over twice as frequently as the remaining three. Since we want to evenly prioritize classifying each type of accident, we turn to balanced accuracy, which is computed as the (unweighted) average recall over all classes, as a more reliable metric for our use case.
IV-C Modality Alignment, Pre-training and Hyperparameters
The core TG model consists of three encoders for three different modalities: text, vision, and graph. For our classification task, a MLP classifier head is placed at the end of the model to output scores for each of the four classes. The text and vision encoders are both kept frozen throughout all of our experiments as we train the SGE to align its embedding space with those of text and vision using Symmetric Cross Entropy Loss Wang et al. (2019). To evaluate the effectiveness of this training regiment, we analyze the classification performance with and without training for modality alignment. We also experiment with the effect of pre-training the SGE before conducting the alignment training. As an additional pre-training baseline, we use a pre-trained model provided by Malawade et al. (2022), which was trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al. (2017). Finally, we perform tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.
IV-D Implementation Details
We train the SGE for 200 epochs on the default hyperparameter settings as specified by Yu et al. (2020). For the SGE pre-training experiments, we select the SGE with the lowest validation loss. We also experiment with the provided SGE pre-trained checkpoint that was trained on synthetic traffic scenes generated by CARLA. For the modality alignment training, we keep most of the default hyperparameter settings, varying batch size (32 to 128) and the number of epochs (8 to 20).
V Results and Discussion
V-A SGE Pre-training Results
As seen in Table I, pre-training the SGE on the DoTA dataset allows it to perform the classification task approximately 13 percentage points better than random classification, which would have an expected accuracy of 25%. This suggests that the spatial and temporal information encoded in the graph embeddings produced by the SGE contain useful information towards accident classification. Unsurprisingly, the model pre-trained on the synthetic CARLA dataset does not achieve as strong performance on our task as the model trained on DoTA. This is understandable, given that the former was trained to classify whether or not a traffic maneuver is deemed to be risky Malawade et al. (2022) rather than to classify traffic accidents. Furthermore, the synthetic traffic scenes in the CARLA dataset used by the authors do not vary as much as those seen in DoTA, and do not contain examples from the four classes used for our task. However, by further training the CARLA model on the DoTA dataset, we are able to achieve similar results to training on DoTA alone (compare two rightmost columns of Table I). For the remaining experiments in the paper, when unspecified, we use the SGE pre-trained only on the DoTA dataset.
TABLE I: SGE Performance with varying pre-train data.
Class-balanced accuracy scores are indicated with a (b).
Pre-training Data CARLA DoTA CARLA+DoTA
Accuracy
Train - 62.72 64.09
Train (b) - 64.78 61.42
Validation - 42.49 41.91
Validation (b) - 40.79 41.42
Test 17.32 37.88 36.03
Test (b) 26.33 38.05 37.81
TABLE II: Test Accuracy Comparison between three system settings. In the first setting, the SGE is not used and the inputs to the system are only from the video and text modalities (the SGG and SGE are unused and stage III-C from the pipeline is skipped). In the second, the SGE is used and aligned with the vision and text encoders before running classification (full architecture is used, but SGE pre-training done in stage III-B may be skipped). In the third, the SGE is used but not aligned with the other encoders (stage III-C is skipped).
Train Setting No SGE SGE with Alignment SGE without Alignment
Pre-training data None None§ CARLA DoTA DoTA
Test accuracy 57.04 54.97 52.66 58.66 57.97
Test accuracy (b) 53.22 53.53 53.97 56.97 57.74
Column label 1 2 3 4 5
V-B SGE with Alignment Results
The primary classification results after training for alignment can be seen in Table II. In Column 1, we verify the baseline accuracy of the model when the scene graph modality is not used at all. Looking at the balanced accuracy score, it is quite interesting to note that the performance under this setting is very close to that of the aligned model when no pre-training data is used or if only the CARLA dataset is used (Columns 2, 3). However, pre-training the model on the DoTA dataset (Column 4) shows a boost in performance by about 3%. This boost suggests that the SG embeddings may be encoding some information about the traffic scenes that cannot be fully captured by either the language or vision modalities.
V-B1 Pre-trained SGE with Alignment
Regarding the case when the model is pre-trained on the DoTA dataset, the model actually has a slightly better performance without undergoing the alignment training (Column 5) versus with alignment (Column 4). This seems to suggest that the model is not learning further useful information during alignment training. One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur Zhang et al. (2023). Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and relies heavily on using a large batch size at each step Radford et al. (2021). This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples. However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training. With a larger batch size (and possibly increasing the number of training epochs), it is possible that the hypothesized trend of higher accuracy when training for alignment would be present. To this end, we show some additional results in Table III, which indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II). It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
TABLE III: Pre-training on DoTA with different configurations. “e” is number of epochs, “bs” is batch size. Column 4 in Table II is the same as Column 1 in this table.
Pre-training on DoTA & SGE with Alignment
10e, bs32 20e, bs32 20e, bs64 10e, bs128
Test accuracy 58.66 58.2 57.51 60.51
Test accuracy (b) 56.97 57.77 57.53 57.72
Column label 1 2 3 4
V-B2 No Pre-trained SGE with Alignment
Another surprising point that arises from the results presented in Table II is that there is very little improvement when training the SGE for alignment from scratch (Column 2) over the baseline “No SGE” case (Column 1). Similar to the case where we pre-trained on DoTA, it is once again possible that further increasing the batch size could be the solution. This is suggested by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table.
No Pre-training & SGE with Alignment
10e, bs32§ 10e, bs64 10e, bs128
Test accuracy 54.97 57.51 59.35
Test accuracy (b) 53.53 53.94 56.00
Column label 1 2 3
VI Conclusion and Future Work
In this work, we have shown that encoding traffic information in the form of a scene graph is beneficial towards the goal of accident classification. This is made clear by the pre-training encoder results which show the SGE’s ability to beat a random classifier at this task. It was further illustrated that the scene graph information can serve to enhance the performance of a vision-language classifier by fusing information from all three modalities. Finally, although we were not able to show an improved score for the classifier after alignment, it should be noted that alignment does not have a negative effect on performance, and we demonstrate that further training and fine-tuning may improve the score beyond the unaligned case.
There are several areas where future work can be done in this task. Firstly, although three modalities are used, little exploration has been conducted on the language modality. In principle, captions for each video should be derived from the video itself, as the SGs are. This would allow for the captions to be fed into the model to fine-tune the classifier head and run inference with a greater signal coming from the language modality. Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right). Enabling a semantic extension of the generated SGs may enhance the signal provided from this modality. Similarly, by modeling relations between all objects in the scene, this pipeline can be expanded for the non-ego case as well, and perhaps can be used on more classes in the DoTA dataset or similar use cases. Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al. (2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components. This work has demonstrated that further fine-tuning and alignment training shows promise to improve results. Experiments with different modality fusion methods as well as classification heads can also lead to improvements.
Appendix A Generated Scene Graph Example
Figure 2: A scene graph generated using the rs2v Malawade et al. [2021] tool. Starting from a video frame (Raw Image), the SGG first detects objects in the scene (Object Detection Image) and generates the BEV (Bird’s Eye Image) before creating the scene graph representation (SceneGraph Image). This scene graph shows the ego car relative to two other vehicles, one categorized as in the left lane, the other right. The closer vehicle is recognized as being near collision (with the edge attribute “near_coll”), whereas the farther vehicle is registered in the scene graph as simply being “visible”.
Appendix B Captions
TABLE V: Two sets of captions that were experimented with for training during alignment. All results shown in this report used Style B. For fine-tuning the classification head, all examples were trained with the same generic caption, “An accident as a result of a vehicle doing something.”
Accident Class Caption Style A Caption Style B
Moving Ahead or Waiting The vehicle is moving ahead or waiting in the accident. An accident as a result of a vehicle moving into another vehicle.
Oncoming The vehicle is hitting an oncoming vehicle in the accident. An accident as a result of a vehicle hitting an oncoming vehicle.
Turning The vehicle is turning in the accident. An accident as a result of a vehicle turning.
Lateral The vehicle is moving laterally in the accident. An accident as a result of a vehicle moving laterally.

### v2

@@
 An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from reoccurring.
-The task of being able to classify a traffic scene as a specific type of accident is the focus of this work.
-We approach the problem by likening a traffic scene to a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges.
-This representation of an accident can be referred to as a scene graph, and is used as input for an accident classifier.
-Better results can be obtained with a classifier that fuses the scene graph input with representations from vision and language.
-This work introduces a multi-stage, multimodal pipeline to pre-process videos of traffic accidents, encode them as scene graphs, and align this representation with vision and language modalities for accident classification.
+This work focuses on classification of traffic scenes into specific accident types.
+We approach the problem by representing a traffic scene as a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges.
+This representation of a traffic scene is referred to as a scene graph, and can be used as input for an accident classifier.
+Better results are obtained with a classifier that fuses the scene graph input with visual and textual representations.
+This work introduces a multi-stage, multimodal pipeline that pre-processes videos of traffic accidents, encodes them as scene graphs, and aligns this representation with vision and language modalities before executing the classification task.
 When trained on 4 classes, our method achieves a balanced accuracy score of 57.77% on an (unbalanced) subset of the popular Detection of Traffic Anomaly (DoTA) benchmark, representing an increase of close to 5 percentage points from the case where scene graph information is not taken into account.
@@
 I Introduction
-The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicle (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022).
+The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicles (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022; Huang et al., 2024; Stoler et al., 2024a, b).
 A major aspect of this task is to efficiently and accurately recognize different types of traffic accidents, with the ultimate goal of preventing them.
-In this work, we approach this challenge by modeling traffic scenes using scene graphs.
+We approach this challenge by modeling traffic scenes using scene graphs.
 To improve classification performance, the graph representation is further aligned with representations from the vision and language modalities within contrastively-trained foundation models.
 Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset Yao et al.
-(2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline to pre-process videos, encode them as scene graphs, and align these representations with vision and language modalities to classify traffic accidents.
-In this paper, we propose Scene-Traffic-Graph Inference(STGi) as a unified system for accident classification.
-First, we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al.
+(2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline that pre-processes videos, encodes them as scene graphs, and aligns these representations with vision and language modalities before classifying traffic accidents.
+We propose STGi as a unified system for accident classification.
+To start, in STGi we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al.
 (2021).
 Then, we use a scene graph encoder (SGE) from the same authors Malawade et al.
-(2022) to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al.
+(2022) of rs2v to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al.
 (2021) and X-CLIP Ni et al.
-(2022), respectively.
-X-CLIP directly expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames.
-In essence, scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and are all used together to classify traffic accident scenes.
+(2022), respectively [^1] CLIP is a multi-modal model that is capable of zero-shot image classification by measuring text-image similarity, and its main components are a text encoder and an image encoder.
+X-CLIP expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames..
+Scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and all are used together to classify traffic accident scenes.
 A summary of contributions is as follows:
-  • We introduce Scene-Traffic-Graph Inference (STGi), a novel method for traffic accident classification which leverages scene graphs to capture the essential features of a traffic accident.
-  • We demonstrate that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
-  • Experiments in this work demonstrate that aligning the scene graph modality with vision and language give similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
+  • We introduce Scene-Traffic-Graph Inference (STGi), a novel classification method that leverages scene graphs to capture essential features from traffic accidents.
+  • We show that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
+  • Experiments in this work demonstrate that aligning the scene graph modality with vision and language gives similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
 Figure 1: An overview of the STGi Architecture.
@@
 Guo et al.
-(2023) extracts similar semantic information from a traffic scene, but the focus is on generating a knowledge graph, accounting for added traffic information present in a scene such as road signs and lanes.
+(2023) extract semantic information from a traffic scene as well, but their focus is largely on generating graphs from road signs and lanes to understand how a traffic scene may be interpreted from the perspective of a driver.
 Qasemi et al.
-(2023) also use a knowledge graph to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents.
-Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a knowledge graph.
-The use of a knowledge graph in this case is helpful in contextualizing the traffic scene in a broader traffic monitoring ontology and for answering general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene.
-Instead, our method is better suited for the task of accident classification by leveraging scene graphs to capture relationships among objects that may be involved in a traffic accident in a given video example.
-Combining the use of scene graphs with knowledge graphs to model traffic settings, Dimasi (2023) makes use of a knowledge graph to boost the performance of the RelTR (Relation Transformer) model Cong et al.
+(2023) use a knowledge graph (KG) to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents.
+Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a KG.
+It is interesting to contrast their KG-based approach, which helps contextualize the traffic scene in a broader traffic monitoring ontology, with our scene graph method, which builds a unique graphical structure from each scene.
+In the former case, the KG can answer general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene.
+Instead, our method is better suited for the task of accident classification by leveraging SGs to capture relationships among objects that may be involved in a traffic accident in a given video example.
+Combining the use of SGs with KGs to model traffic settings, Dimasi (2023) makes use of a KG to boost the performance of the RelTR (Relation Transformer) model Cong et al.
 (2022), an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
 II-B Leveraging Scene Graphs with Neural Networks
-There are many possible designs and applications for leveraging the versatile structure of scene graphs to enhance the performance of neural networks Li et al.
+There are many possible designs and applications for leveraging the versatile structure of SGs to enhance the performance of neural networks Li et al.
 (2019); Khademi and Schulte (2020); Rana et al.
+(2023); Saxena et al.
 (2024).
-More generally, Zhang et al.
-(2019) describe the ability of scene graphs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features.
-Although their study is not specific to our traffic domain, they claim that scene graphs derived from images can capture the essential visual features and may outperform images for VQA tasks.
-Many works have focused on automatically generating scene graphs that can easily be encoded into neural networks for various downstream tasks.
-The rs2v tool Malawade et al.
-(2021) generates scene graphs for individual frames of traffic videos and is applied to our work.
+Zhang et al.
+(2019) describe the ability of SGs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features.
+Although their study is not specific to our traffic domain, they claim that SGs derived from images can capture the essential visual features and may outperform images for VQA tasks.
+Many works have focused on automatically generating SGs that can easily be encoded into neural networks for various downstream tasks.
+For example, the rs2v tool Malawade et al.
+(2021) generates SGs for individual frames of traffic videos and is applied to our work.
 Alternatively, Nag et al.
-(2023) design a tool specifically for modeling videos with scene graphs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers.
-However, they train on the Action Genome Ji et al.
-(2019) dataset and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
+(2023) design a tool specifically for modeling videos with SGs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers.
+However, they train on the Action Genome dataset Ji et al.
+(2019) and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
 II-C Multimodal Alignment with Contrastive Representations
 Radford et al.
-(2021) proposed a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model.
-From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as aligning audio Wu et al.
-(2022) and haptic data Tatiya et al.
+(2021) propose a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model.
+From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as audio Wu et al.
+(2022) and haptic modalities Tatiya et al.
 (2024).
 Similar to our implementation, Koch et al.
-(2023) use CLIP’s text encoder to contrastively train 3D scene graphs.
+(2023) use CLIP’s text encoder to contrastively train 3D scene graphs, which are graphs that implicitly contain spatial information of objects in a scene, and can also model other relationships between these objects Koch et al.
+(2023).
 The work by Huang et al.
@@
 III Methodology
-To approach the problem of traffic accident classification, a multi-stage pipeline was developed.
-The pipeline, which is represented in Figure 1, can be summarized in four stages and is described in the following sections:
+To approach classification of traffic accidents, a multi-stage pipeline was developed.
+The pipeline, which is represented in Figure 1, can be summarized in four stages that are described in the following sections:
   1. 1.
@@
 Some do work for our needs — in particular, the roadscene2vec (rs2v) tool created by Malawade et al.
-(2021) which we leverage for generating scene graphs from traffic video frames.
-To use rs2v, we first sample a fixed amount of frames from each video, and text captions are generated manually for each class.
+(2021) that we leverage for generating scene graphs from traffic video frames.
+To use rs2v, we first sample a fixed amount of frames from each video, and for each class we manually generate text captions.
 Then, feeding in a series of frames from a traffic scene, the rs2v generator uses an object detector and keeps only the relevant entities relating to traffic (such as road, cars, pedestrians, etc.), filtering out other detected objects.
-Next, after generating a “bird’s eye view” (BEV) projection of the image and approximating the relative location of each object in this projection, edges are connected between nearby entities in the scene, and vehicles are mapped to lanes.
+Next, after generating a “bird’s eye view” (BEV) projection of the image and approximating the relative location of each object in this projection, edges are added to connect nearby entities in the scene and to map vehicles to lanes.
 This forms the scene graph for a specific frame from a video sample.
-The SGG is an important component of this pipeline as it defines the elements in the scene graph modality.
-An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used to create edges of varying attributes, such as very near or visible, relating a pair of objects).
+The SGG is a crucial component of this pipeline as it defines the elements in the scene graph modality.
+An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used when creating edges labeled with varying attributes, such as very near or visible, relating a pair of objects).
 The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene Malawade et al.
 (2022).
-In our case, we iteratively sample and adjust the BEV parameters and proximity thresholds based on the output quality of the SGs produced for various scenes in DoTA.
+In our case, we generate SGs for a random subset of DoTA traffic scenes, visually inspect them for errors (e.g., the SGG labels the edge relating the ego vehicle to a vehicle directly in front of it as visible rather than very near), and then manually adjust the BEV parameters and proximity thresholds in an iterative manner.
 We do this to select one configuration to generate the SGs for our task, although it remains an inherent challenge to generalize the parameter settings for the entire dataset.
 Some examples of the generated SGs can be seen in Appendix A.
-For text, we manually compose captions describing each of the four accident classes (see Appendix B), pairing each caption with videos from its respective class to form the training examples.
+For text, we manually compose captions describing each of the four accident classes (see Appendix B), and pair each caption with videos from its respective class to form the training examples.
 We also experiment with two sets of captions, however, we focus the scope of this work primarily on the scene graph generation process from video frames rather than caption generation.
 It should be noted that although these captions are used during training for aligning the SGE, they are not used during inference on the final model because they are not from our dataset and provide a one-to-one mapping directly to the accident classes.
-Through fine-tuning the classification head using these captions, the model can achieve 100% accuracy this way.
-Instead, we fine-tune and test the classification head using the same caption as noted in Appendix B for all examples.
+Through fine-tuning the classification head using these captions, the model can achieve 100% accuracy.
+Instead, we fine-tune and test the classification head using the same caption, as noted in Appendix B, for all examples.
 III-B Scene Graph Encoder
-In stage two, the SGE encodes scene graphs to fixed-length embeddings.
+In stage two, the SGE encodes sequences of scene graphs derived from a given video to fixed-length embeddings.
 To do this, a multirelational graph convolutional network (MRGCN) as employed by Malawade et al.
@@
 IV-A Dataset
-Given the nature of our multimodal approach, we require data in the form of videos, text, and scene graphs.
-Regarding the videos, we focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al.
-(2022) for this task.
-DoTA is a dataset consisting of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place.
-For our use case of this dataset, question one is irrelevant since we only sample frames from the anomalous portion of the video based on the annotations of the DoTA authors.
-Additionally, question two is less essential for us since we use a SGG, which is meant to extract the necessary objects from video frames and model their relevant spatial relationships for our task.
-We therefore prioritize providing an answer for the third question, that is, classifying different traffic anomalies.
-However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to only a subset of four of the available classes.
-In addition to the provided video classes, Yao et al.
-(2022) also divides the videos into two types, ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego.
-Every video class contains videos of both types.
+Our multimodal approach requires data in the form of videos, text, and scene graphs.
+We focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al.
+(2022) for videos for this task.
+DoTA is a dataset of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place.
+For our use case, question one is irrelevant since we only sample frames from the anomalous portion of the video based on the annotations of the DoTA authors.
+Question two is also less essential for us since we use a SGG, which extracts the necessary objects from video frames and models their relevant spatial relationships.
+We therefore prioritize answering the third question, that is, classifying different traffic anomalies.
+However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to a subset of four of the available classes.
+In addition to the video classes, Yao et al.
+(2022) also divide the videos into two types (which are different from classes), ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego.
+Every video class has videos of both types (e.g., there are both ego and non-ego videos that fall into the turning class).
 By default, rs2v is designed to only model relationships between the ego vehicle and other objects, not among other objects themselves, so we further restrict our task to the set of ego videos available in DoTA.
-In total, this leaves us with 4 classes and 2,163 videos.
-The classes are labeled as follows, describing the type movement of the vehicle as the accident occurs: moving_ahead_or_waiting, oncoming, turning, and lateral.
+In total, this leaves us with four classes and 2,163 videos.
+The classes labels, describing the vehicle movement as the accident occurs, are as follows: moving_ahead_or_waiting, oncoming, turning, and lateral.
 IV-B Baselines and Metrics
-In order to evaluate the usefulness of representing traffic scenes graphically to classify them, we first verify whether a SGE alone is superior to randomly guessing an accident class.
-We later evaluate the effectiveness of using SGs in a multimodal classifier by comparing the difference in the performance between using just text and video embeddings versus adding in the signal from the SG modality.
-For this and other results, the primary metrics we focus on are accuracy and balanced accuracy.
-Accuracy is simply defined as the proportion of the correctly classified predictions out of the total number of predictions and remains an interpretable metric that is relatively simple to compare across different benchmarks.
-There is, however, significant imbalance in the dataset, with the turning class appearing over twice as frequently as the remaining three.
-Since we want to evenly prioritize classifying each type of accident, we turn to balanced accuracy, which is computed as the (unweighted) average recall over all classes, as a more reliable metric for our use case.
-IV-C Modality Alignment, Pre-training and Hyperparameters
-The core TG model consists of three encoders for three different modalities: text, vision, and graph.
+In order to evaluate the usefulness of representing traffic scenes graphically for classification, we verify whether a SGE alone is superior to randomly guessing an accident class.
+We then assess the effectiveness of SGs in a multimodal classifier by comparing the difference in performance between using just text and video embeddings versus including the SG modality.
+Regarding metrics, we focus on accuracy and balanced accuracy.
+Accuracy is simply defined as the proportion of the correctly classified predictions out of the total number of predictions and is an interpretable metric that can be easily compared across benchmarks.
+However, there is significant imbalance in the dataset, with the turning class appearing over twice as frequently as the other three.
+We use balanced accuracy, computed as the (unweighted) average recall over all classes, to give equal priority to each accident class.
+IV-C Modality Alignment, Pre-training, and Hyperparameters
+The core STGi model consists of three encoders for three different modalities: text, vision, and graph.
 For our classification task, a MLP classifier head is placed at the end of the model to output scores for each of the four classes.
@@
 We also experiment with the effect of pre-training the SGE before conducting the alignment training.
-As an additional pre-training baseline, we use a pre-trained model provided by Malawade et al.
-(2022), which was trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al.
+As an additional pre-training baseline, we use a model provided by Malawade et al.
+(2022), which was pre-trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al.
 (2017).
-Finally, we perform tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.
+Finally, we conduct tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.
 IV-D Implementation Details
@@
 (2020).
-For the SGE pre-training experiments, we select the SGE with the lowest validation loss.
-We also experiment with the provided SGE pre-trained checkpoint that was trained on synthetic traffic scenes generated by CARLA.
-For the modality alignment training, we keep most of the default hyperparameter settings, varying batch size (32 to 128) and the number of epochs (8 to 20).
+After the SGE pre-training experiments, we select the SGE with the lowest validation loss to use for the modality alignment training.
+During alignment training, we keep most of the default hyperparameter settings, except for varying the batch size (32 to 128) and the number of epochs (8 to 20).
 V Results and Discussion
 V-A SGE Pre-training Results
-As seen in Table I, pre-training the SGE on the DoTA dataset allows it to perform the classification task approximately 13 percentage points better than random classification, which would have an expected accuracy of 25%.
-This suggests that the spatial and temporal information encoded in the graph embeddings produced by the SGE contain useful information towards accident classification.
+As seen in Table I, pre-training the SGE on DoTA allows it to perform classification approximately 13 percentage points better than random classification, which would have an expected accuracy of 25%.
+This suggests that spatial and temporal information encoded in the graph embeddings produced by the SGE contain useful information towards accident classification.
 Unsurprisingly, the model pre-trained on the synthetic CARLA dataset does not achieve as strong performance on our task as the model trained on DoTA.
@@
 (2022) rather than to classify traffic accidents.
-Furthermore, the synthetic traffic scenes in the CARLA dataset used by the authors do not vary as much as those seen in DoTA, and do not contain examples from the four classes used for our task.
-However, by further training the CARLA model on the DoTA dataset, we are able to achieve similar results to training on DoTA alone (compare two rightmost columns of Table I).
+Furthermore, the scenes in CARLA do not vary as much as those in DoTA, and do not contain examples from the four classes we use.
+However, by further training the CARLA model on DoTA, we achieve similar results to training on DoTA alone (compare two rightmost columns of Table I).
 For the remaining experiments in the paper, when unspecified, we use the SGE pre-trained only on the DoTA dataset.
@@
 Pre-training Data CARLA DoTA CARLA+DoTA
-Accuracy
-Train - 62.72 64.09
-Train (b) - 64.78 61.42
-Validation - 42.49 41.91
-Validation (b) - 40.79 41.42
-Test 17.32 37.88 36.03
-Test (b) 26.33 38.05 37.81
+Train accuracy - 62.72 64.09
+Train accuracy (b) - 64.78 61.42
+Validation accuracy - 42.49 41.91
+Validation accuracy (b) - 40.79 41.42
+Test accuracy 17.32 37.88 36.03
+Test accuracy (b) 26.33 38.05 37.81
 TABLE II: Test Accuracy Comparison between three system settings.
@@
 In the third, the SGE is used but not aligned with the other encoders (stage III-C is skipped).
+All values are obtained using a batch size of 32 and 10 epochs of training.
 Train Setting No SGE SGE with Alignment SGE without Alignment
@@
 Test accuracy (b) 53.22 53.53 53.97 56.97 57.74
-Column label 1 2 3 4 5
 V-B SGE with Alignment Results
@@
 In Column 1, we verify the baseline accuracy of the model when the scene graph modality is not used at all.
-Looking at the balanced accuracy score, it is quite interesting to note that the performance under this setting is very close to that of the aligned model when no pre-training data is used or if only the CARLA dataset is used (Columns 2, 3).
-However, pre-training the model on the DoTA dataset (Column 4) shows a boost in performance by about 3%.
+Looking at the balanced accuracy score, it is interesting to note that the performance on this setting is only slightly under that of the aligned model when no pre-training data is used or if only the CARLA dataset is used (Columns 2, 3).
+However, pre-training the model on the DoTA dataset (Column 4) shows a further boost in performance by about 3%.
 This boost suggests that the SG embeddings may be encoding some information about the traffic scenes that cannot be fully captured by either the language or vision modalities.
@@
 (2023).
-Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and relies heavily on using a large batch size at each step Radford et al.
+Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step Radford et al.
 (2021).
@@
 However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training.
-With a larger batch size (and possibly increasing the number of training epochs), it is possible that the hypothesized trend of higher accuracy when training for alignment would be present.
-To this end, we show some additional results in Table III, which indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II).
+We hypothesize that a larger batch size and/or number of training epochs would increase the accuracy gain of the training alignment step.
+To test this hypothesis, we carried out some additional experiments, the results of which are shown in Table III, and indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II).
 It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
-TABLE III: Pre-training on DoTA with different configurations. “e” is number of epochs, “bs” is batch size. Column 4 in Table II is the same as Column 1 in this table.
+TABLE III: Alignment training with different hyperparameter configurations, using a SGE pre-trained on DoTA.
+“e” is number of epochs, “bs” is batch size for the alignment training.
+Column 4 in Table II is the same as Column 1 in this table.
 Pre-training on DoTA & SGE with Alignment
@@
 Test accuracy (b) 56.97 57.77 57.53 57.72
-Column label 1 2 3 4
 V-B2 No Pre-trained SGE with Alignment
-Another surprising point that arises from the results presented in Table II is that there is very little improvement when training the SGE for alignment from scratch (Column 2) over the baseline “No SGE” case (Column 1).
+Another surprising point arising from Table II, which was mentioned at the start of Section V-B, is that there is very little improvement when training the SGE for alignment from scratch (Column 2) over the baseline “No SGE” case (Column 1).
 Similar to the case where we pre-trained on DoTA, it is once again possible that further increasing the batch size could be the solution.
-This is suggested by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
+This is corroborated by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
 TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table.
@@
 Test accuracy (b) 53.53 53.94 56.00
-Column label 1 2 3
 VI Conclusion and Future Work
@@
 This is made clear by the pre-training encoder results which show the SGE’s ability to beat a random classifier at this task.
-It was further illustrated that the scene graph information can serve to enhance the performance of a vision-language classifier by fusing information from all three modalities.
-Finally, although we were not able to show an improved score for the classifier after alignment, it should be noted that alignment does not have a negative effect on performance, and we demonstrate that further training and fine-tuning may improve the score beyond the unaligned case.
+It was further illustrated that scene graph information can enhance the performance of a vision-language classifier by fusing information from all three modalities.
+Finally, although we were not able to show an improved score for the classifier after alignment, it should be noted that alignment training using an appropriate hyperparameter configuration does not have a negative effect on performance, and we demonstrated that further training and fine-tuning may improve the score beyond the unaligned case.
 There are several areas where future work can be done in this task.
@@
 Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right).
-Enabling a semantic extension of the generated SGs may enhance the signal provided from this modality.
-Similarly, by modeling relations between all objects in the scene, this pipeline can be expanded for the non-ego case as well, and perhaps can be used on more classes in the DoTA dataset or similar use cases.
+Enabling a semantic extension of the generated SGs may enhance the signal obtained from this modality.
+Similarly, by modeling relations between all objects in the scene, this pipeline could be expanded for the non-ego case as well, and perhaps could be used on more accident classes in the DoTA dataset or similar use cases.
+Regarding the dataset, although we attempt to resolve the data imbalance by focusing on the balanced accuracy metric, it would be interesting to explore data augmentation techniques as well.
 Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al.
 (2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components.
-This work has demonstrated that further fine-tuning and alignment training shows promise to improve results.
-Experiments with different modality fusion methods as well as classification heads can also lead to improvements.
-Appendix A Generated Scene Graph Example
+Given the modular design of our system, another fascinating avenue of exploration would be to incorporate an additional modality (e.g., audio) to include in the alignment step before fusing the embeddings and fine tuning the classification head.
+Experiments with different modality fusion methods as well as classification heads may also lead to improvements.
+Appendix A Generated Scene Graph Examples
 Figure 2: A scene graph generated using the rs2v Malawade et al.
@@
 The closer vehicle is recognized as being near collision (with the edge attribute “near_coll”), whereas the farther vehicle is registered in the scene graph as simply being “visible”.
+Figure 3: Starting from a video frame, the SGG first detects objects in the scene and generates the BEV image before creating the scene graph representation.
+This scene graph shows the ego car relative to a person on a scooter, categorized as in the left lane.
+The person is recognized as being near collision, with the edge attribute “near_coll”.
+Note that there is a vehicle detected in the distance (right part of detection image, highlighted in blue) but it is not included in the graph since it is too far.
 Appendix B Captions

### v3

@@
 Recognizing a traffic accident is an essential part of any autonomous driving or road monitoring system.
-An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from reoccurring.
-This work focuses on classification of traffic scenes into specific accident types.
+An accident can appear in a wide variety of forms, and understanding what type of accident is taking place may be useful to prevent it from recurring.
+This work focuses on classifying traffic scenes into specific accident types.
 We approach the problem by representing a traffic scene as a graph, where objects such as cars can be represented as nodes, and relative distances and directions between them as edges.
@@
 Index Terms:
-Neuro-symbolism, Vision-Language Models, Foundation Models, Autonomous Driving, Perception
+Neuro-symbolic Reasoning, Vision-Language Models, Multimodal Learning, Autonomous Driving
 I Introduction
-The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicles (AV) and road infrastructure systems (Xu et al., 2021; Yao et al., 2022; Qasemi et al., 2023; Francis et al., 2022; Huang et al., 2024; Stoler et al., 2024a, b).
+The task of understanding traffic scenarios is one with ever-increasing importance, particularly for the advancement of Autonomous Vehicles (AV) and road infrastructure systems [1, 2, 3, 4, 5, 6, 7].
 A major aspect of this task is to efficiently and accurately recognize different types of traffic accidents, with the ultimate goal of preventing them.
@@
 To improve classification performance, the graph representation is further aligned with representations from the vision and language modalities within contrastively-trained foundation models.
-Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset Yao et al.
-(2022), we build Scene-Traffic-Graph Inference (STGi), a multi-stage, multimodal pipeline that pre-processes videos, encodes them as scene graphs, and aligns these representations with vision and language modalities before classifying traffic accidents.
-We propose STGi as a unified system for accident classification.
-To start, in STGi we generate scene graphs using the roadscene2vec (rs2v) tool Malawade et al.
-(2021).
-Then, we use a scene graph encoder (SGE) from the same authors Malawade et al.
-(2022) of rs2v to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP Radford et al.
-(2021) and X-CLIP Ni et al.
-(2022), respectively [^1] CLIP is a multi-modal model that is capable of zero-shot image classification by measuring text-image similarity, and its main components are a text encoder and an image encoder.
+Starting with video clips of traffic incidents, procured from the Detection of Traffic Anomaly (DoTA) dataset [2], we build Traffic-Scene-Graph Inference (TSGi), a multi-stage, multimodal pipeline that pre-processes videos, encodes them as scene graphs, and aligns these representations with vision and language modalities before classifying traffic accidents.
+We propose TSGi as a unified system for accident classification.
+To start, in TSGi we generate scene graphs using the roadscene2vec (rs2v) tool [8].
+Then, we use a scene graph encoder (SGE) from the same authors [9] of rs2v to obtain encodings that are aligned to encodings of textual and visual inputs, obtained from CLIP [10] and X-CLIP [11], respectively [^1] CLIP is a multi-modal model that is capable of zero-shot image classification by measuring text-image similarity, and its main components are a text encoder and an image encoder.
 X-CLIP expands upon CLIP’s image encoder to include an attention mechanism to model inter-frame communication and to generate a new embedding representation from video frames..
-Scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and all are used together to classify traffic accident scenes.
+In essence, scene graphs are treated as a new modality (or ‘view’) that is aligned with text and video signals, and are all used together to classify traffic accident scenes.
 A summary of contributions is as follows:
-  • We introduce Scene-Traffic-Graph Inference (STGi), a novel classification method that leverages scene graphs to capture essential features from traffic accidents.
+  • We introduce Traffic-Scene-Graph Inference (TSGi), a novel method for traffic accident classification that leverages scene graphs to capture essential features from traffic accidents.
   • We show that the added signal from a scene graph modality can enhance the performance of a video-language traffic accident classifier by nearly 5 percentage points.
   • Experiments in this work demonstrate that aligning the scene graph modality with vision and language gives similar results to omitting alignment training, although increasing the batch size and training time during alignment shows a trend of increasing scores and the potential in further improving classification results.
-Figure 1: An overview of the STGi Architecture.
+Figure 1: An overview of the TSGi Architecture.
 Beginning with video and text inputs, video frames are sampled and used to generate scene graphs.
@@
 Several works have taken the approach of modeling traffic scenarios using scene graphs (SGs).
-Yu et al.
-(2020) propose one such representation that forms the basis of the SGs produced in our work.
-They define a graph structure with nodes representing entities such as vehicles and pedestrians while edges represent relations between the objects and are labeled by distance category (e.g., visible, very near) and orientation (e.g., in front of, beside).
-Zipfl and Zöllner (2021) also model these objects in a traffic scene, further accounting for relative speeds of objects in the graph edges.
-Guo et al.
-(2023) extract semantic information from a traffic scene as well, but their focus is largely on generating graphs from road signs and lanes to understand how a traffic scene may be interpreted from the perspective of a driver.
-Qasemi et al.
-(2023) use a knowledge graph (KG) to enhance a video-language model for the related task of Question Answering in the domain of traffic accidents.
-Their method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a KG.
-It is interesting to contrast their KG-based approach, which helps contextualize the traffic scene in a broader traffic monitoring ontology, with our scene graph method, which builds a unique graphical structure from each scene.
-In the former case, the KG can answer general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene.
-Instead, our method is better suited for the task of accident classification by leveraging SGs to capture relationships among objects that may be involved in a traffic accident in a given video example.
-Combining the use of SGs with KGs to model traffic settings, Dimasi (2023) makes use of a KG to boost the performance of the RelTR (Relation Transformer) model Cong et al.
-(2022), an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
+One such representation proposed in [12] forms the basis of the SGs produced in our work.
+The authors define a graph structure with nodes representing entities such as vehicles and pedestrians while edges represent relations between the objects and are labeled by distance category (e.g., visible, very near) and orientation (e.g., in front of, beside).
+Objects in a traffic scene are also modeled in [13], further accounting for relative speeds of objects in the graph edges.
+In [14], semantic information is extracted from a traffic scene as well, but the focus is largely on generating graphs from road signs and lanes to understand how a traffic scene may be interpreted from the perspective of a driver.
+For the related task of Question Answering in the domain of traffic accidents, [3] uses a knowledge graph to enhance a video-language model.
+The method is to inject domain-specific knowledge by first auto-generating video annotations and storing semantic information about a traffic scene in a knowledge graph.
+It is interesting to contrast such an approach, which uses a knowledge graph to help contextualize a traffic scene in a broader traffic monitoring ontology, with our scene graph method, which builds a unique graphical structure from each scene.
+In the former case, the knowledge graph can answer general traffic accident-related questions, such as the impact of having more vehicles in a traffic scene, but is less effective in identifying accident types based on the interactions among vehicles in a specific traffic scene.
+Instead, our method is better suited for the task of accident classification by leveraging scene graphs to capture relationships among objects that may be involved in a traffic accident in a given video example.
+Combining the use of scene graphs with knowledge graphs to model traffic settings, [15] makes use of a knowledge graph to boost the performance of the RelTR (Relation Transformer) model [16], an encoder-decoder model used for scene graph generation by framing it as a set prediction problem.
 II-B Leveraging Scene Graphs with Neural Networks
-There are many possible designs and applications for leveraging the versatile structure of SGs to enhance the performance of neural networks Li et al.
-(2019); Khademi and Schulte (2020); Rana et al.
-(2023); Saxena et al.
-(2024).
-Zhang et al.
-(2019) describe the ability of SGs to enhance Visual Question Answering (VQA) systems as an alternative to solely focusing on vision and language features.
-Although their study is not specific to our traffic domain, they claim that SGs derived from images can capture the essential visual features and may outperform images for VQA tasks.
-Many works have focused on automatically generating SGs that can easily be encoded into neural networks for various downstream tasks.
-For example, the rs2v tool Malawade et al.
-(2021) generates SGs for individual frames of traffic videos and is applied to our work.
-Alternatively, Nag et al.
-(2023) design a tool specifically for modeling videos with SGs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers.
-However, they train on the Action Genome dataset Ji et al.
-(2019) and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
+There are many possible designs and applications for leveraging the versatile structure of scene graphs to enhance the performance of neural networks [17, 18, 19, 20].
+As discussed in [21], Visual Question Answering (VQA) systems may be enhanced using scene graphs as an alternative to solely focusing on vision and language features.
+Although their study is not specific to our traffic domain, [21] claims that scene graphs derived from images can capture the essential visual features and may outperform images for VQA tasks.
+Many works have focused on automatically generating scene graphs that can easily be encoded into neural networks for various downstream tasks.
+For example, the rs2v tool [8] generates scene graphs for individual frames of traffic videos and is applied to our work.
+Alternatively, [22] designs a tool specifically for modeling videos with scene graphs which accounts for spatial and temporal relationships between frames based on sequence modeling with transformers.
+However, they train on the Action Genome dataset [23] and produce graphs that are unrelated to traffic scenarios, making this tool less applicable to our work.
 II-C Multimodal Alignment with Contrastive Representations
-Radford et al.
-(2021) propose a model that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model.
-From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as audio Wu et al.
-(2022) and haptic modalities Tatiya et al.
-(2024).
-Similar to our implementation, Koch et al.
-(2023) use CLIP’s text encoder to contrastively train 3D scene graphs, which are graphs that implicitly contain spatial information of objects in a scene, and can also model other relationships between these objects Koch et al.
-(2023).
-The work by Huang et al.
-(2023) aims to enhance CLIP’s performance by injecting scene graph information into the embedding generated by CLIP’s text encoder before contrastive learning is performed between the image and text embedding spaces.
+In [10], a model is proposed that uses a contrastive training strategy to align text captions with images for the Contrastive Language–Image Pre-training (CLIP) model.
+From this, several works have focused on aligning encoders of additional modalities to those trained for CLIP, such as audio [24] and haptic modalities [25].
+Similar to our implementation, [26] uses CLIP’s text encoder to contrastively train 3D scene graphs, which are graphs that implicitly contain spatial information of objects in a scene, and can also model other relationships between these objects [26].
+The work in [27] aims to enhance CLIP’s performance by injecting scene graph information into the embedding generated by CLIP’s text encoder before contrastive learning is performed between the image and text embedding spaces.
 To the best of our knowledge, however, our work is the first to treat scene graphs as an additional modality that is grounded to a text and video encoder.
-There are also many possible strategies for fusing the outputs of a multimodal model before the downstream task at hand, such as early and late stage fusion involving concatenation Sleeman et al.
-(2022); Pawłowski et al.
-(2023), merging Sleeman et al.
-(2022), or sampling from a shared embedding space Gao et al.
-(2020).
+There are also many possible strategies for fusing the outputs of a multimodal model before the downstream task at hand, such as early and late stage fusion involving concatenation [28, 29], merging [28], or sampling from a shared embedding space [30].
 We experiment with several of these late fusion techniques on the three modality embeddings after training for alignment.
@@
 Various scene graph generators (SGGs) are available for this, although not many fit our specific requirements for a generator that is capable of identifying and encoding features unique to traffic accidents, such as different types of vehicle collisions.
-Some do work for our needs — in particular, the roadscene2vec (rs2v) tool created by Malawade et al.
-(2021) that we leverage for generating scene graphs from traffic video frames.
+Some do work for our needs — in particular, the roadscene2vec (rs2v) tool [8] that we leverage for generating scene graphs from traffic video frames.
 To use rs2v, we first sample a fixed amount of frames from each video, and for each class we manually generate text captions.
@@
 An important step before employing this tool is to calibrate the BEV and adjust the proximity thresholds (which are used when creating edges labeled with varying attributes, such as very near or visible, relating a pair of objects).
-The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene Malawade et al.
-(2022).
+The authors specifically note the challenges they faced when adjusting these settings with regard to the DoTA dataset, and mention that for best results, they needed to calibrate their model for each traffic scene [9].
 In our case, we generate SGs for a random subset of DoTA traffic scenes, visually inspect them for errors (e.g., the SGG labels the edge relating the ego vehicle to a vehicle directly in front of it as visible rather than very near), and then manually adjust the BEV parameters and proximity thresholds in an iterative manner.
@@
 In stage two, the SGE encodes sequences of scene graphs derived from a given video to fixed-length embeddings.
-To do this, a multirelational graph convolutional network (MRGCN) as employed by Malawade et al.
-(2022) is used, which includes an attention mechanism along with LSTMs to model the spatial and temporal relations of the scene graphs generated for a given video.
+To do this, a multirelational graph convolutional network (MRGCN) as employed in [9] is used, which includes an attention mechanism along with LSTMs to model the spatial and temporal relations of the scene graphs generated for a given video.
 The SGE may be pre-trained for the classification task before aligning the scene graphs with the language and vision modalities.
 III-C Multimodal Alignment
-For stage three, we use the CLIP model Radford et al.
-(2021), but replace the CLIP image encoder with X-CLIP Ni et al.
-(2022), a pre-trained video encoder, to accept videos rather than images.
+For stage three, we use the CLIP model [10], but replace the CLIP image encoder with X-CLIP [11], a pre-trained video encoder, to accept videos rather than images.
 We leave CLIP’s text encoder in place.
@@
 This requires the selection of a method to fuse the signal from the three modalities followed by training a small network.
-For this, we experiment with different late fusion techniques such as taking a weighted linear combination of the modality outputs Kaliciak et al.
-(2014) and training various MLP classifiers Kiela et al.
-(2018) with and without activations on top of the concatenated embeddings.
-Based on these experiments, we choose the approach of concatenating the vectors from the three modalities as described by Sleeman et al.
-(2022) and, following Wu et al.
-(2022), train a 2-layer MLP with ReLU activations as our classification head.
+For this, we experiment with different late fusion techniques such as taking a weighted linear combination of the modality outputs [31] and training various MLP classifiers [32] with and without activations on top of the concatenated embeddings.
+Based on these experiments, we choose the approach of concatenating the vectors from the three modalities as described in [28] and, following [24], train a 2-layer MLP with ReLU activations as our classification head.
 IV Experimental Design
@@
 Our multimodal approach requires data in the form of videos, text, and scene graphs.
-We focus on using the Detection of Traffic Anomaly (DoTA) dataset Yao et al.
-(2022) for videos for this task.
+We focus on using the Detection of Traffic Anomaly (DoTA) dataset [2] for videos for this task.
 DoTA is a dataset of 4,677 curated videos of traffic anomalies from YouTube, and was specifically designed to answer the following three questions about each video: (1) When does the anomalous event (i.e., accident) start and end; (2) Where are the anomalous regions of each video frame; and (3) What type of anomaly is taking place.
@@
 However, given the multitude of classes and the inherent challenge in distinguishing them even with the naked eye, we limit the problem to a subset of four of the available classes.
-In addition to the video classes, Yao et al.
-(2022) also divide the videos into two types (which are different from classes), ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego.
+In addition to the video classes, [2] also divides the videos into two types (which are different from classes), ego (where the vehicle from which the video recording is captured is involved in the accident), and its opposite, non-ego.
 Every video class has videos of both types (e.g., there are both ego and non-ego videos that fall into the turning class).
 By default, rs2v is designed to only model relationships between the ego vehicle and other objects, not among other objects themselves, so we further restrict our task to the set of ego videos available in DoTA.
-In total, this leaves us with four classes and 2,163 videos.
+In total, this leaves us with 4 classes and 2,163 videos.
 The classes labels, describing the vehicle movement as the accident occurs, are as follows: moving_ahead_or_waiting, oncoming, turning, and lateral.
@@
 IV-C Modality Alignment, Pre-training, and Hyperparameters
-The core STGi model consists of three encoders for three different modalities: text, vision, and graph.
+The core TSGi model consists of three encoders for three different modalities: text, vision, and graph.
 For our classification task, a MLP classifier head is placed at the end of the model to output scores for each of the four classes.
-The text and vision encoders are both kept frozen throughout all of our experiments as we train the SGE to align its embedding space with those of text and vision using Symmetric Cross Entropy Loss Wang et al.
-(2019).
+The text and vision encoders are both kept frozen throughout all of our experiments as we train the SGE to align its embedding space with those of text and vision using Symmetric Cross Entropy Loss [33].
 To evaluate the effectiveness of this training regiment, we analyze the classification performance with and without training for modality alignment.
 We also experiment with the effect of pre-training the SGE before conducting the alignment training.
-As an additional pre-training baseline, we use a model provided by Malawade et al.
-(2022), which was pre-trained on a synthetic traffic dataset generated using the CARLA traffic simulator Dosovitskiy et al.
-(2017).
+As an additional pre-training baseline, we use a model provided in [9], which was pre-trained on a synthetic traffic dataset generated using the CARLA traffic simulator [34].
 Finally, we conduct tests to evaluate the performance of different hyperparameter settings (batch size, training time) for the alignment training.
 IV-D Implementation Details
-We train the SGE for 200 epochs on the default hyperparameter settings as specified by Yu et al.
-(2020).
+We train the SGE for 200 epochs on the default hyperparameter settings as specified in [12].
 After the SGE pre-training experiments, we select the SGE with the lowest validation loss to use for the modality alignment training.
@@
 Unsurprisingly, the model pre-trained on the synthetic CARLA dataset does not achieve as strong performance on our task as the model trained on DoTA.
-This is understandable, given that the former was trained to classify whether or not a traffic maneuver is deemed to be risky Malawade et al.
-(2022) rather than to classify traffic accidents.
+This is understandable, given that the former was trained to classify whether or not a traffic maneuver is deemed to be risky [9] rather than to classify traffic accidents.
 Furthermore, the scenes in CARLA do not vary as much as those in DoTA, and do not contain examples from the four classes we use.
@@
 TABLE I: SGE Performance with varying pre-train data.
-Class-balanced accuracy scores are indicated with a (b).
+Class-balanced accuracy scores are indicated with a (b)
 Pre-training Data CARLA DoTA CARLA+DoTA
@@
 In the third, the SGE is used but not aligned with the other encoders (stage III-C is skipped).
-All values are obtained using a batch size of 32 and 10 epochs of training.
+All values are obtained using a batch size of 32 and 10 epochs of training
 Train Setting No SGE SGE with Alignment SGE without Alignment
@@
 This seems to suggest that the model is not learning further useful information during alignment training.
-One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur Zhang et al.
-(2023).
-Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step Radford et al.
-(2021).
+One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur [35].
+Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step [10].
 This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples.
@@
 “e” is number of epochs, “bs” is batch size for the alignment training.
-Column 4 in Table II is the same as Column 1 in this table.
+Column 4 in Table II is the same as Column 1 in this table
 Pre-training on DoTA & SGE with Alignment
@@
 This is corroborated by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
-TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table.
+TABLE IV: Training for alignment from scratch with different configurations. Column 2 in Table II is the same as Column 1 in this table
 No Pre-training & SGE with Alignment
@@
 Regarding the dataset, although we attempt to resolve the data imbalance by focusing on the balanced accuracy metric, it would be interesting to explore data augmentation techniques as well.
-Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al.
-(2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components.
+Additionally, the MRGCN-based architecture of the SGE was taken directly from [12], but perhaps it can be further improved by modifying either its spatial or temporal modeling components.
 Given the modular design of our system, another fascinating avenue of exploration would be to incorporate an additional modality (e.g., audio) to include in the alignment step before fusing the embeddings and fine tuning the classification head.
@@
 Appendix A Generated Scene Graph Examples
-Figure 2: A scene graph generated using the rs2v Malawade et al.
-[2021] tool.
+Figure 2: A scene graph generated using the rs2v [8] tool.
 Starting from a video frame (Raw Image), the SGG first detects objects in the scene (Object Detection Image) and generates the BEV (Bird’s Eye Image) before creating the scene graph representation (SceneGraph Image).